OpenStudy (anonymous):

I Have a Question, I Know the Answer, But i Wanna Know Why It's The Answer, Can Someone Explain?

OpenStudy (anonymous):

As part of putting the finishing touches on her time machine, Misty has to solve this equation. 1 5 (x+7)−8=27 She completes the problem as seen in the steps below. The time machine won't work! Her assistant looks at the problem and tells her the solution is x=168.

OpenStudy (anonymous):

Sorry here

OpenStudy (anonymous):

Its Step 2.

OpenStudy (anonymous):

(1/5)(x+7) -8 = 27 (1/5)x+1.4-8 = 27 (1/5)x = 33.6 x = 168